Given that sinu + sinq - (27/65) and tan[(u + q)/ 2] = (3/7), find cos[(u - q)/ 2]
1
Expert's answer
2021-02-16T07:56:37-0500

Solution


"\qquad\qquad
\begin{aligned}
\sin u+\sin q&= \frac{27}{65}\\
2\sin (\frac{u+q}{2}).\cos(\frac{u-q}{2})&= \small \frac{27}{65}........................(1)\\
\\
\sin(u+q)&=2\sin (\frac{u+q}{2}).\cos(\frac{u+q}{2}).........................(2)\\
\\
\sin(u+q)&= \frac{2\tan(\frac{u+q}{2})}{1-\tan^2(\frac{u+q}{2})}\\
&= \frac{2\times\frac{3}{7}}{1-\frac{9}{49}}\\
&= \frac{21}{20}.................................(3)\\
\\
(2)=(3)\\
\\
2\sin(\frac{u+q}{2}).\cos(\frac{u+q}{2})&= \frac{21}{20}...............................(4)\\
\\
\tan(\frac{u+q}{2})&= \frac{3}{7}\\
\frac{\sin(\frac{u+q}{2})}{\cos(\frac{u+q}{2})}&= \frac{3}{7}\implies \cos(\frac{u+q}{2})=\frac{7}{3}\sin(\frac{u+q}{2})............(5)\\
\\
Substituting \,(5)\,in\,(4),
\\
2\sin(\frac{u+q}{2}).\frac{7}{3}.\sin(\frac{u+q}{2})&=\frac{21}{20}\\
\sin^2(\frac{u+q}{2})&=\frac{9}{40}\\
\sin(\frac{u+q}{2})&=\pm\frac{3}{2\sqrt{10}}\\
&=+0.474\,or\,-0.474\\
\\
According\,to\,(1),\\
When\,\,\,\sin(\frac{u+q}{2})=+0.474&\implies\cos(\frac{u-q}{2})=+0.438\\
When\,\,\,\sin(\frac{u+q}{2})=-0.474 &\implies\cos(\frac{u-q}{2})=-0.438
\end{aligned}"